OpenStudy (a1234):

(16a)/(sqrt 6a^3) The answer key says (8 sqrt 6a)/(3a), but I can't see how.

OpenStudy (anonymous):

\[\frac{ 16a }{ \sqrt{6a^3} }\] This is the equation?

OpenStudy (a1234):

Yes

OpenStudy (a1234):

@MrsDavidRandle Can you help me?

OpenStudy (anonymous):

I don't think so, sorry.

OpenStudy (anonymous):

You have to use the rule for dividing exponents. But division is just like multiplication, except it's with reciprocal. So check this out. You know that x^a * x^b = x^(a+b), yeah? If you don't, then plug in a number into x, a, and b. For example 3, 4, and 5. 3^4 * 3^5 = 3*3*3*3 * 3*3*3*3*3. But there's eight 3's in there, so that equals 3^8. Anyway, what if b is negative? x^a * x^(-b) = x^a / x^b. You divide them because x^(-1) = 1/x. It's the reciprocal. So since you add exponents when you multiply, you subtract them when you divide.

OpenStudy (anonymous):

Now, back to your problem. \[\frac{16a}{\sqrt{6a^3}}\] Remember that the sqrt is actually an exponent of 1/2. Also (x^a)^b = x^(a*b). So the denominator is actually sqrt(6) * a^(3/2), since (a^3)^(1/2) = a^(3/2), right? Now, we're dividing a / a^(3/2). The numerator a has an exponent of 1, so 1-3/2 = -1/2. So, a / a^(3/2) = a^(-1/2) = 1 / √a. But it needs to be "rationalized," so that's actually √a / a. If we rationalize the 1 / √6, we get √6 / 6. Let's rewrite what we have so far: \[\frac{16 \sqrt{6} \sqrt{a} }{ 6 a}\] Now we simply divide 16/6 to get 8/3.

OpenStudy (kropot72):

Multiplying numerator and denominator by \[\sqrt{6a ^{3}}\] gives \[\frac{16a \sqrt{6a ^{3}}}{6a ^{3}}\] Dividing numerator and denominator by 2a^2 gives \[\frac{\frac{8a}{a ^{2}}\sqrt{6a ^{3}}}{3a}\] Now take a^2 outside the radical sign, making it a: \[\frac{\frac{8a ^{2}}{a ^{2}}\sqrt{6a}}{3a}\] Cancelling the a^2 gives the answer.

OpenStudy (a1234):

Thanks TurtleMuffin and Kropot! Kropot, your method seems kind of different from the one my textbook shows my, but it works.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Oh right, we could have rationalized from the get-go. The beauty of math is that one can manipulate any expression in any kind of way, as long as you follow algebraic rules!